OpenStudy (anonymous):

what is 8 devide by 4 1/2 will give medal for best answers

OpenStudy (whpalmer4):

Dividing by a fraction can be accomplished by multiplying by the reciprocal of the fraction. You make the reciprocal by inverting the fraction. For example, \[\frac{3}{\frac{1}{2}} = 3*\frac{2}{1} = \frac{6}{1} = 6\]

OpenStudy (anonymous):

0.47058823529 RadDeg x! Inv sin ln π cos log e tan √ Ans EXP xy ( ) % AC 7 8 9 ÷ 4 5 6 × 1 2 3 - 0 . = +

OpenStudy (anonymous):

lol what tf is that @sel95

OpenStudy (whpalmer4):

Convert \(4\frac{1}{2}\) into an improper fraction, and then invert it and multiply it by \(8\) to divide \(8\) by \(4\frac{1}{2}\)

OpenStudy (anonymous):

This Convert 412 into an improper fraction, and then invert it and multiply it by 8 to divide 8 by 412. Done wrong

OpenStudy (whpalmer4):

@sel95 what do you mean, it is done wrong? \[4\frac{1}{2} = \frac{4*2+1}{2} = \frac{9}{2}\]\[\frac{8}{4\frac{1}{2}} = 8*\frac{2}{9} = \frac{8*2}{9} = \frac{16}{9}\] \[\frac{16}{9} = 1.777777...\]\[8/4.5 = 1.777777...\]

OpenStudy (anonymous):

@whpalmer4 I did it wrong I believe